# Differentially Private Synthetic Data via Foundation Model APIs 1: Images

### v2

@@ 3             $S_{t+1}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
             $histogram_{t}^{c}\leftarrow{\textsf{DP\_NN\_HISTOGRAM}}{}\left(private\_samples^{c},S_{t}^{c},\sigma,H\right)$   // See Alg. 2
3
@@
             $\mathcal{P}_{t}\leftarrow histogram_{t}^{c}/\mathrm{sum}(histogram_{t}^{c})$   // $\mathcal{P}_{t}$ is a distribution on $S_{t}^{c}$
             $S_{t}^{\prime}\leftarrow$ draw $N_{\mathrm{syn}}/|C|$ samples with replacement from $\mathcal{P}_{t}$   // $S_{t}^{\prime}$ is a multiset
@@ 5             $S_{t+1}^{c}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
6
$S_{\mathrm{syn}}\leftarrow S_{\mathrm{syn}}\cup\left\{(x,c)|x\in S_{T}^{c},c\in C\right\}$ return $S_{\mathrm{syn}}$
@@
Appendix D Theoretical Evidence for Convergence of PE
@@
Convergence of Non-Private Evolution. We first analyze Alg. 1 when no noise is added to the histograms (i.e., we set $\sigma=0$ and $H=0$ in Alg. 1). We show that in this case, the evolution algorithm does converge to the private distribution in $O(d)$ iterations where $d$ is the dimension of the embedding space. Under some reasonable modeling assumptions (see App. E), we prove the following theorem. Here $D$ is the diameter of $S_{\mathrm{priv}}$, $L\approx$ number of variations of each point in $S_{t}^{\prime}$ that are added to $S_{t+1}$ in Alg. 1 of Alg. 1.
@@
Convergence of Private Evolution. To get some intuition on the working of PE in the presence of noise, we make a simplifying assumption. We will assume that there are $B$ identical copies of each private point in $S_{\mathrm{priv}}$. We call $B$ as multiplicity. Note that for any DP algorithm to converge to $S_{\mathrm{priv}}$, we need that the private data is well clustered with some minimum cluster size. Any cluster with too few points cannot be represented in $S_{\mathrm{syn}}$, because that would violate DP. And when there is a cluster of $B$ private points and the generated points in $S_{t}$ are still far from this cluster, then it is likely that all the $B$ private points will have a common closest point in $S_{t}$ [^10] In fact, it is easy to formally prove that there will be a common approximate nearest neighbor., i.e., they all vote for the same point in $S_{t}$ as a single entity. Therefore multiplicity is a reasonable modeling assumption to make to understand the working of PE. Note that, actually it is very easy to find $S_{\mathrm{priv}}$ exactly using DP Set Union (Gopi et al., 2020) with the multiplicity assumption. The point of Thm. 2 is to give intuition about why PE works in practice; it is proved in § E.2 under the same assumptions as in Thm. 1.
@@
Appendix E Proofs of PE Convergence Theorems
@@
Appendix F Intrinsic Dimension of Image Embeddings
@@
    We use VARIATION_API from App. I to obtain 3000 image variations of $x$: $x_{1},\ldots,x_{3000}$, and their corresponding inception embeddings $g_{1},...,g_{3000}\in{}^{2048}$. 3000 is chosen so that the number of variations is larger than the embedding dimension.
@@
We conduct the above process with the variation degree $[98,96,94,92,90,88,86,84,82,80,78,76,74,72,70,68,66,64,62,60]$ utilized in the CIFAR10 experiments (see App. I). We additionally add a variation degree of 100 which is the highest variation degree in the API that was used to generate the initial samples. We plot the estimated intrinsic dimension v.s. variation degree in Fig. 9. The raw original singular values of $M/\sqrt{3000}$ for variation degree=60 are in Fig. 10 (other variation degrees have similar trend). Two key observations are:
@@
Appendix G Relation of DPWA to prior work
@@ G.1 Why Not Just Use DP Clustering?
    It exploits the intrinsic dimension of the manifold of images in the embedding space which is much smaller than the embedding dimension (see Thm. 2 and App. F) and
@@
Appendix H Implementation Details on Label Condition
@@
Conditional pre-trained networks/APIs. We first explain our implementation when the pre-trained networks or APIs support conditional inputs such as class labels or text prompts. When generating the initial population using RANDOM_API (Alg. 1), we will either randomly draw labels from all possible labels when no prior public information is available (which is what we do in CIFAR10 and Camelyon17 experiments where we randomly draw from all possible ImageNet classes), or use the public information as condition input (e.g., the text prompt used in Stable Diffusion experiments; see App. K). In the subsequent VARIATION_API calls (Alg. 1), for each image, we will use its associated class label or text prompt as the condition information to the API, and the output samples from VARIATION_API will be associated with the same class label or text prompt as the input sample. For example, if we use an image with “peacock” class to generate variations, all output images will be associated with “peacock” class for future VARIATION_API calls. Note that throughout the above process, all the condition inputs to the pre-trained networks/APIs are public information; they have nothing to do with the private classes.
@@
Appendix I More Details and Results on CIFAR10 Experiments
@@
API implementation. RANDOM_API follows the standard diffusion model sampling process. VARIATION_API is implemented with SDEdit (Meng et al., 2021), which adds noise to input images and lets the diffusion model denoise them. We use DDIM sampler (Song et al., 2020) and the default noise schedule to draw samples. Note that these choices are not optimal; our results can potentially be improved by using better noise schedules and the full DDPM sampling (Ho et al., 2020) which are known to work better. The implementation of the above APIs is straightforward without touching the core modeling part of diffusion models and is similar to the standard API implementations in Stable Diffusion (App. K).
Hyperparameters. We set number of iterations $T=20$, lookahead degree $k=8$, and number of generated samples $N_{\mathrm{syn}}=50000$. For RANDOM_API and VARIATION_API, we use DDIM sampler with 100 steps. For VARIATION_API, we use SDEEdit (Meng et al., 2021) by adding noise till $[98,96,94,92,90,88,86,84,82,80,78,76,74,72,70,68,66,64,62,60]$ timesteps for each iteration respectively. These timesteps can be regarded as the $v$ parameter in § 3.3.
@@
For the experiments in Fig. 5, we use noise multiplier $\sigma=\sqrt{2}$ and threshold $H=2$ (i.e., $t=1$).
@@
Appendix J More Details and Resutls on Camelyon17 Experiments
@@
About the experiments in Fig. 22. For RANDOM_API and VARIATION_API, we use DDIM sampler with 10 steps. For VARIATION_API, we take a 2-stage approach. the first stage, we use DDIM sampler with 10 steps and use SDEEdit (Meng et al., 2021) by adding noise till $[10,10,10,10,9,9,9,9,9,8,8,8,8,8,7,7,7,7]$ timesteps for each iteration respectively. In the second stage, we use DDIM sampler with 40 steps and use SDEEdit (Meng et al., 2021) by adding noise till $[20,19,18,17,16,15,14,13,12,11,10,9,8,7,6,5]$ timesteps for each iteration respectively. These timesteps can be regarded as the $v$ parameter in § 3.3. We use noise multiplier $\sigma=2\cdot\sqrt{2}$ and threshold $H=4$.
@@
Inspescting the ImageNet class labels of the generated samples. As discussed in App. H, each generated samples have an associated ImageNet label. Here, we inspect those labels to understand how PE works in this dataset. The list of the labels and their associated number of generated images are: “honeycomb” (164647), “velvet” (83999), “nematode, nematode worm, roundworm” (35045), “handkerchief, hankie, hanky, hankey” (14495), “bib” (3102), “head cabbage” (934), “bubble” (142), “stole” (72). We see that 54.4% images are with the label “honeycomb”. Many honeycomb images in ImageNet have a “net-like” structure, which shares some similarities with Camelyon17. However, the details, colors, and structures are still different. We compute the FID between honeycomb images and CIFAR10. The FID score is 162.86, which is much higher than the FID score of the final generated images we get (10.66 in Fig. 22). These results show that the existence of the honeycomb class in ImageNet helps with PE, but PE does more than just picking this class.
@@
Appendix K More Details and Results on Stable Diffusion Experiments
@@
Appendix L CINIC Experiments
@@
Appendix M More Ablation Studies
@@
The results are shown in Table 2, which suggests that larger $N_{\mathrm{priv}}$ helps the performance of PE, similar to the observation in DP-SGD (Anil et al., 2021). We hypothesize that the reason is as follows. When we set $N_{\mathrm{syn}}/N_{\mathrm{priv}}=1$, although the signal-noise ratio in the DP Nearest Neighbors Histogram remains constant, larger $N_{syn}$ does allow the generated samples to explore larger space, and therefore it is more likely to get a sample closer to the private data, which helps the convergence of PE. Our theorem in App. D also shows that increasing the number of variations (which is controlled by $N_{syn}$) speeds up the convergence.
@@
Appendix N Generating More Samples
@@
Appendix O Computational Cost Evaluation
@@
In the above evaluation, we did not consider the time of generating more samples beyond $N_{\mathrm{syn}}=50000$, and we discuss its effect here. To generate more samples beyond $N_{\mathrm{syn}}=50000$ samples, the DP fine-tuning method can directly use the fine-tuned model to generate more samples without more fine-tuning. PE can use the approach in § 4.2, where we pass the generated $N_{\mathrm{syn}}$ samples through VARIATION_API to generate more samples (what we did in § 5.1.3; see results thereof). Note that, for the same size of the model, the same number of generated samples, the same diffusion sampler, and the same total number of denoising steps, the running time of PE to generate more samples is smaller than the DP fine-tuning baseline. The reason is as follows. Let’s say we use a diffusion sampler with 100 steps in total. In the DP fine-tuning baseline, generated samples must take all 100 steps (i.e., starting from Gaussian noise, passing the image through the diffusion model 100 times iteratively to get the final generated samples). For PE, VARIATION_API is implemented by SDEdit (Meng et al., 2021) (see App. I), where we add noise to input images and let the diffusion model denoise them starting from the middle of the diffusion process (e.g., adding Gaussian noise to the image, then passing the image through the diffusion model starting from 20th denoising step for all the rest 80 steps iteratively to get the final generated samples). In other words, each generated does not need to go through all 100 steps, but only a fraction of the steps (80 steps in the above example).

### v3

@@ 1 $S_{\mathrm{syn}}\leftarrow\emptyset$
             $histogram_{t}^{c}\leftarrow{\textsf{DP\_NN\_HISTOGRAM}}{}\left(private\_samples^{c},S_{t}^{c},\sigma,H\right)$
              // See Alg. 2
6
@@
             $\mathcal{P}_{t}\leftarrow histogram_{t}^{c}/\mathrm{sum}(histogram_{t}^{c})$
              // $\mathcal{P}_{t}$ is a distribution on $S_{t}^{c}$
             $S_{t}^{\prime}\leftarrow$ draw $N_{\mathrm{syn}}/|C|$ samples with replacement from $\mathcal{P}_{t}$
              // $S_{t}^{\prime}$ is a multiset
@@ 8             $S_{t+1}^{c}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
9
[…]
return $S_{\mathrm{syn}}$
@@
Appendix E Theoretical Evidence for Convergence of PE
@@
Convergence of Non-Private Evolution. We first analyze Alg. 1 when no noise is added to the histograms (i.e., we set $\sigma=0$ and $H=0$ in Alg. 1). We show that in this case, the evolution algorithm does converge to the private distribution in $O(d)$ iterations where $d$ is the dimension of the embedding space. Under some reasonable modeling assumptions (see App. F), we prove the following theorem. Here $D$ is the diameter of $S_{\mathrm{priv}}$, $L\approx$ number of variations of each point in $S_{t}^{\prime}$ that are added to $S_{t+1}$ in Alg. 1 of Alg. 1.
@@
Convergence of Private Evolution. To get some intuition on the working of PE in the presence of noise, we make a simplifying assumption. We will assume that there are $B$ identical copies of each private point in $S_{\mathrm{priv}}$. We call $B$ as multiplicity. Note that for any DP algorithm to converge to $S_{\mathrm{priv}}$, we need that the private data is well clustered with some minimum cluster size. Any cluster with too few points cannot be represented in $S_{\mathrm{syn}}$, because that would violate DP. And when there is a cluster of $B$ private points and the generated points in $S_{t}$ are still far from this cluster, then it is likely that all the $B$ private points will have a common closest point in $S_{t}$ [^10] In fact, it is easy to formally prove that there will be a common approximate nearest neighbor., i.e., they all vote for the same point in $S_{t}$ as a single entity. Therefore multiplicity is a reasonable modeling assumption to make to understand the working of PE. Note that, actually it is very easy to find $S_{\mathrm{priv}}$ exactly using DP Set Union (Gopi et al., 2020) with the multiplicity assumption. The point of Thm. 2 is to give intuition about why PE works in practice; it is proved in § F.2 under the same assumptions as in Thm. 1.
@@
Appendix F Proofs of PE Convergence Theorems
@@
Appendix G Intrinsic Dimension of Image Embeddings
@@
    We use VARIATION_API from App. J to obtain 3000 image variations of $x$: $x_{1},\ldots,x_{3000}$, and their corresponding inception embeddings $g_{1},...,g_{3000}\in{}^{2048}$. 3000 is chosen so that the number of variations is larger than the embedding dimension.
@@
We conduct the above process with the variation degree $[98,96,94,92,90,88,86,84,82,80,78,76,74,72,70,68,66,64,62,60]$ utilized in the CIFAR10 experiments (see App. J). We additionally add a variation degree of 100 which is the highest variation degree in the API that was used to generate the initial samples. We plot the estimated intrinsic dimension v.s. variation degree in Fig. 9. The raw original singular values of $M/\sqrt{3000}$ for variation degree=60 are in Fig. 10 (other variation degrees have similar trend). Two key observations are:
@@
Appendix H Relation of DPWA to prior work
@@ H.1 Why Not Just Use DP Clustering?
    It exploits the intrinsic dimension of the manifold of images in the embedding space which is much smaller than the embedding dimension (see Thm. 2 and App. G) and
@@
Appendix I Implementation Details on Label Condition
@@
Conditional pre-trained networks/APIs. We first explain our implementation when the pre-trained networks or APIs support conditional inputs such as class labels or text prompts. When generating the initial population using RANDOM_API (Alg. 1), we will either randomly draw labels from all possible labels when no prior public information is available (which is what we do in CIFAR10 and Camelyon17 experiments where we randomly draw from all possible ImageNet classes), or use the public information as condition input (e.g., the text prompt used in Stable Diffusion experiments; see App. L). In the subsequent VARIATION_API calls (Alg. 1), for each image, we will use its associated class label or text prompt as the condition information to the API, and the output samples from VARIATION_API will be associated with the same class label or text prompt as the input sample. For example, if we use an image with “peacock” class to generate variations, all output images will be associated with “peacock” class for future VARIATION_API calls. Note that throughout the above process, all the condition inputs to the pre-trained networks/APIs are public information; they have nothing to do with the private classes.
@@
Appendix J More Details and Results on CIFAR10 Experiments
@@
API implementation. RANDOM_API follows the standard diffusion model sampling process. VARIATION_API is implemented with SDEdit (Meng et al., 2021), which adds noise to input images and lets the diffusion model denoise them. We use DDIM sampler (Song et al., 2020) and the default noise schedule to draw samples. Note that these choices are not optimal; our results can potentially be improved by using better noise schedules and the full DDPM sampling (Ho et al., 2020) which are known to work better. The implementation of the above APIs is straightforward without touching the core modeling part of diffusion models and is similar to the standard API implementations in Stable Diffusion (App. L).
Hyperparameters. We set the maximum number of iterations $T=20$, lookahead degree $k=8$, and number of generated samples $N_{\mathrm{syn}}=50000$. For RANDOM_API and VARIATION_API, we use DDIM sampler with 100 steps. For VARIATION_API, we use SDEEdit (Meng et al., 2021) by adding noise till $[98,96,94,92,90,88,86,84,82,80,78,76,74,72,70,68,66,64,62,60]$ timesteps for each iteration respectively. These timesteps can be regarded as the $v$ parameter in § 3.3.
@@
For the experiments in Fig. 5, we use noise multiplier $\sigma=2\sqrt{2}$, threshold $H=4$ (i.e., $t=2$), and $T=5$.
@@
Appendix K More Details and Resutls on Camelyon17 Experiments
@@
About the experiments in Fig. 22. For RANDOM_API and VARIATION_API, we use DDIM sampler with 10 steps. For VARIATION_API, we take a 2-stage approach. the first stage, we use DDIM sampler with 10 steps and use SDEEdit (Meng et al., 2021) by adding noise till $[10,10,10,10,9,9,9,9,9,8,8,8,8,8,7,7,7,7]$ timesteps for each iteration respectively. In the second stage, we use DDIM sampler with 40 steps and use SDEEdit (Meng et al., 2021) by adding noise till $[20,19,18,17,16,15,14,13,12,11,10]$ timesteps for each iteration respectively. These timesteps can be regarded as the $v$ parameter in § 3.3. We use noise multiplier $\sigma=2\cdot\sqrt{2}$ and threshold $H=4$.
@@
Inspescting the ImageNet class labels of the generated samples. As discussed in App. I, each generated samples have an associated ImageNet label. Here, we inspect those labels to understand how PE works in this dataset. The list of the labels and their associated number of generated images are: “honeycomb” (164647), “velvet” (83999), “nematode, nematode worm, roundworm” (35045), “handkerchief, hankie, hanky, hankey” (14495), “bib” (3102), “head cabbage” (934), “bubble” (142), “stole” (72). We see that 54.4% images are with the label “honeycomb”. Many honeycomb images in ImageNet have a “net-like” structure, which shares some similarities with Camelyon17. However, the details, colors, and structures are still different. We compute the FID between honeycomb images and CIFAR10. The FID score is 162.86, which is much higher than the FID score of the final generated images we get (10.66 in Fig. 22). These results show that the existence of the honeycomb class in ImageNet helps with PE, but PE does more than just picking this class.
@@
Appendix L More Details and Results on Stable Diffusion Experiments
@@
Appendix M CINIC Experiments
@@
Appendix N More Ablation Studies
@@
The results are shown in Table 2, which suggests that larger $N_{\mathrm{priv}}$ helps the performance of PE, similar to the observation in DP-SGD (Anil et al., 2021). We hypothesize that the reason is as follows. When we set $N_{\mathrm{syn}}/N_{\mathrm{priv}}=1$, although the signal-noise ratio in the DP Nearest Neighbors Histogram remains constant, larger $N_{syn}$ does allow the generated samples to explore larger space, and therefore it is more likely to get a sample closer to the private data, which helps the convergence of PE. Our theorem in App. E also shows that increasing the number of variations (which is controlled by $N_{syn}$) speeds up the convergence.
@@
Appendix O Generating More Samples
@@
Appendix P Computational Cost Evaluation
@@
In the above evaluation, we did not consider the time of generating more samples beyond $N_{\mathrm{syn}}=50000$, and we discuss its effect here. To generate more samples beyond $N_{\mathrm{syn}}=50000$ samples, the DP fine-tuning method can directly use the fine-tuned model to generate more samples without more fine-tuning. PE can use the approach in § 4.2, where we pass the generated $N_{\mathrm{syn}}$ samples through VARIATION_API to generate more samples (what we did in § 5.1.3; see results thereof). Note that, for the same size of the model, the same number of generated samples, the same diffusion sampler, and the same total number of denoising steps, the running time of PE to generate more samples is smaller than the DP fine-tuning baseline. The reason is as follows. Let’s say we use a diffusion sampler with 100 steps in total. In the DP fine-tuning baseline, generated samples must take all 100 steps (i.e., starting from Gaussian noise, passing the image through the diffusion model 100 times iteratively to get the final generated samples). For PE, VARIATION_API is implemented by SDEdit (Meng et al., 2021) (see App. J), where we add noise to input images and let the diffusion model denoise them starting from the middle of the diffusion process (e.g., adding Gaussian noise to the image, then passing the image through the diffusion model starting from 20th denoising step for all the rest 80 steps iteratively to get the final generated samples). In other words, each generated does not need to go through all 100 steps, but only a fraction of the steps (80 steps in the above example).